I sent these before but for some reason i could not open it could you please help again. I have three problems to do.

it say use linear combinations to solve the system of linear.

question 1. x + 2y = 5
3x - 2y =7

question 2. x+y = 1
2x -3y =12

question 3.x - y = -4
x +2y = 5

please help

Answer by venugopalramana(3286)

SEE THIS AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK
you gave me a site to go on and i can't get on it i am trying to see how to do these because i am out of school right now and hve no one to show me
use linear combinations to solve the system linera equation
THESE ARE ALL SIMILAR PROBLEMS.I SHALL SHOW ONE AND YOU MAY TRY THE OTHERS.
1. x + 2 =5....................EQN.I
3x - 2y =7.....................EQN.II
1.THERE ARE 2 UNKNOWNS AND 2 EQNS.
2.IF THERE IS 1 UNKNOWN AND 1 EQN.WE CAN SOLVE IT
3.SO WE TRY TO REMOVE ONE UNKNOWN BY ELIMINATION.
EQN.I GIVES X=5-2=3
EQN.II ..3*3-2Y=7
2Y=9-7=2
Y=1
2. x + y = 1......................III
2x -3y = 12........................IV
EQN.I*2-EQN.II
2X+2Y-2X+3Y=2-12=-10
5Y=-10
Y=-2
PUT THIS IN EQN.III
X-2=1
X=3....ETC..
3. x -y = -4
x + 2y = 5